OpenStudy (anonymous):

Find the length of side x of the triangle, Round too nearest hundredth. Pic in question.

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Its simple. Just use Pythagoras theorem. 10^2 + 24^2 = x^2 Solve for x and that'll be the answer.

OpenStudy (anonymous):

As it is right angled triangle: \[Hypotenuse^2 = Base^2 + Height^2\] \[H = \sqrt{B^2 + h^2}\] \[H = \sqrt{10^2 + 24^2} = \sqrt{676}\] What is the square root of 676?? That will be your answer..

OpenStudy (anonymous):

ah I see so it's 26, thanks

OpenStudy (anonymous):

Welcome dear..